find a number such that,one-fourth of a number is 30 more than 7
Answer by stanbon(75887)

find a number such that,one-fourth of a number is 30 more than 7
---
Equation:
(1/4)x = 7+30
(1/4)x = 37
x = 4*37
x = 148
===============
Cheers,
Stan H.